OpenStudy (anonymous):

Which of these points lie on the line that passes through (6,2) with slope m=2/3? Select all answers that are correct. A) (3,0) B) (-4,0) C) (3,-2) D) (-3,-4)

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

\[y-y_a=m(x-x_a) \\ m=\frac{2}{3} \\ (x_a,y_a)=(6,2)\]

OpenStudy (anonymous):

plug in the values to the first equation to make the equation you are looking for. Then you can test the points you are given by plugging them into the equation and seeing if they are true or not

OpenStudy (anonymous):

Can you show me how to do that? I don't really understand @romero plzzz

OpenStudy (anonymous):

You don't understand how to get the equation first?

OpenStudy (anonymous):

no i need help plugging in points to check which ones are my answers @Romero

OpenStudy (anonymous):

Just plug them into to your equation. You do have your equation right? Anyways both sides need to be the same in order to be true.

OpenStudy (anonymous):

You can't have 2=4 or 5=9

OpenStudy (anonymous):

Ok so that eliminates which?

OpenStudy (anonymous):

if the point lies in the line then you should have soemthing like this 6=6

OpenStudy (anonymous):

I don't know you have to test them out by yourself.

OpenStudy (anonymous):

Ok can u at least tell me how many of the options there are?

OpenStudy (anonymous):

What do you mean by that.? If you have the equation you should be able to do this without a problem. Write down and tell me the equation of the line.